# Thermal conductivity minimum with thickness in ultrathin films

### v2

@@ Abstract
Theoretical frameworks based on the Boltzmann equation, including scattering of phonons that bump off the boundaries of the film, as described by the Fuchs-Sondheimer (FS) theory Fuchs (1938); Sondheimer (1952), remain invaluable to explain the increase of thermal conductivity with increasing film thickness in the range of tens to hundreds of nanometers McGaughey et al. (2011). However, recent numerical investigations have revealed new unexpected behaviors upon further shrinking the thickness, that cannot be described by traditional theories. In particular, molecular simulations have highlighted the presence of a minimum in the thermal conductivity of ultrathin silicon films at about 1-2 nm thickness Fu et al. (2020). A similar behaviour, although without a clear minimum and much shallower, has been reported for argon films Fu et al. (2020). As of today, current theoretical framework cannot explain these observation of the minimum in thermal conductivity at 1-2 nm thickness in silicon films and the flattening behaviour in argon films.
Here we apply the nanoconfinement description of phonons in ultrathin few-layer films based on recent developments Yu et al. (2022); Zaccone (2024) and derive a new fundamental framework for the thermal conductivity of ultrathin films that fully takes into account the restriction in the reciprocal space occupancy of phonons due to confinement Balandin and Wang (1998). The new theory is verified in excellent semi-quantitative comparison with recent molecular dynamics (MD) simulations data in terms of the predicted thickness dependence.
In particular, the origin of the minimum is traced back to a crossover in the phonon density of states from a confinement-induced cubic law in frequency to the quadratic Debye law. This effect, in turn, originates from a redistribution of phonon momentum states in reciprocal space associated with a transition from the trivial Debye-sphere homotopy group to a non-trivial manifold with $\mathbb{Z}$ homotopy group.
@@
[…]
where $C_{v}(\omega)$ is the frequency-dependent volumetric heat capacity, $v(\omega)$ is the phonon group velocity, and $\tau(\omega)$ is the phonon scattering time. The latter receives contribution from various processes, that include: Rayleigh-type scattering due to isolated defects, anharmonic phonon-phonon scattering, boundary scattering etc. As we are going to specialize on a simple monoatomic system such as silicon, we can use the original Debye-Peierls model without accounting for optical phonons explicitly. Then, the spectral heat capacity is given by the energy of a mode with frequency $\omega$: $U=g(\omega)\langle n(\omega)\rangle\hbar\omega$, where $g(\omega)$ is the phonon density of states (DOS), and $\langle n(\omega)\rangle=\frac{\hbar\omega}{e^{\hbar\omega/k_{B}T}-1}$ is the Bose-Einstein occupation factor. Taking the derivative of the internal energy of mode $\omega$ with respect to temperature $T$, and using the Debye DOS Kittel (1996): $g(\omega)=\frac{3V}{2\pi^{2}}\frac{\omega^{2}}{v^{3}}$, the spectral heat capacity per unit volume is then:
@@
Finally, the frequency-dependent mean free path of the phonons, which is the inverse of the average relaxation time, is given by Ward and Broido (2010); Walker and Pohl (1963); Braun et al. (2016); Callaway (1959):
$\Lambda_{bulk}(\omega)\equiv\tau^{-1}(\omega)=\frac{v}{L}+A\omega^{4}+B\omega^{2}e^{-C/T}$ (3)
where the first term on the r.h.s. is a boundary scattering term McGaughey et al. (2011); Larkin and McGaughey (2014); Braun et al. (2016) which suffices to describe size-effects in thin films with thickness $L>10$ nm. The second term on the r.h.s. represents the Rayleigh-type scattering of phonons by diluted defects and impurities (a similar quartic form is found also for more generic atomic-scale disorder Baggioli and Zaccone (2022)). The last term on the r.h.s. is the anharmonic contribution from inelastic three-phonon (Umklapp) scattering, famously theorized by Peierls Peierls (1955) and subsequently computed by Klemens Klemens and Simon (1951).
@@
Figure 1: Left: diametral section of the corresponding geometry of $k$-space, where the Debye sphere (of radius $k_{D}$, in red) contains two symmetric ”hollow” spheres (hole pockets of radius $\pi/L$, in white) of forbidden states. These are phonon states in $k$-space that remain unoccupied due to confinement along the $z$-axis of the thin film in real space. Upon further decreasing the film thickness $L$, eventually the Debye spherical surface gets distorted, as shown in the right figure. At this point the phonon momentum states in reciprocal space are strongly redistributed, with a spreading out of states on the distorted Debye surface and many states being pushed towards the core of the red figure, resulting in a faster growth of the number of states with $k$, hence with $\omega$, explaining the higher, cubic, exponent, of the DOS derived in Yu et al. (2022).
However, additional effects arise when the film thickness $L$ drops below 10 nm, which are not taken into account by the above theory framework. These new effects are brought about by the spatial confinement of the quasiparticles, and can affect both thermal Zaccone (2024) and electronic transport Zaccone (2025a), as well as the superconductivity Travaglino and Zaccone (2023); Ummarino and Zaccone (2024). In a nutshell, phonons described as quantum waves have a certain wavelength. Along the confined ($z$) direction of the film, only wavelengths smaller than the thickness $L$ are allowed to exist and to populate the Debye sphere in reciprocal space. On the other hand, in the unconfined $x-y$ plane, phonons with any wavelength are allowed. Along a generic direction identified by the angle $\theta$ (measured with respect to the confinement axis $z$), only phonons with wavelength:
@@
are allowed to populate the Debye sphere in $k$-space Zaccone (2025b). Phonons with larger $\lambda$ are forbidden and the corresponding points in $k$-space form two symmetric hole-pocket spheres of prohibited states within the Debye sphere, cf. Fig. 1 in Zaccone (2025b) and Figs. 1-2 in Phillips et al. (2021). As long as these two hole-pocket spheres are within the surface of the Debye sphere, the DOS is still provided by the Debye DOS quoted above. However, as the thickness $L$ is decreased, the two hole-pocket ”hollow” spheres grow up to the point that they come out of the Debye sphere (right figure in Fig. 1). At this point, the DOS is no longer given by the Debye DOS but can still be derived in simple analytical form, as follows.
@@
which shows a cubic dependence on frequency, $\sim\omega^{3}$, consistent with both experimental results (from inelastic neutron scattering) and molecular dynamics simulations reported in Ref. Yu et al. (2022). Notably, this $\omega^{3}$ behavior is observed in both crystalline and fully amorphous thin films. Furthermore, this $\omega^{3}$ law is approximately observed in MD simulations of ultrathin silicon films and ultrathin argon films Turney et al. (2010); Fu et al. (2020). This cubic DOS for ultrathin films leads to a characteristic $\sim T^{4}$ of the heat capacity with temperature at low temperatures (instead of Debye’s $\sim T^{3}$), which appears in agreement with the available data Zaccone (2024).
@@
This physical model thus explains how the confinement leads to a relative increase in the density of low-frequency vibrational modes, as indeed observed in simulations and experiments Yu et al. (2022); Turney et al. (2010); Fu et al. (2020). In particular, the excess of vibrational modes over the Debye prediction is controlled by the crossover in Eq. (10) occurring at $\omega_{\times}=\frac{2\pi}{L}v$, itself a function of the film thickness $L$. Clearly, as $L$ decreases, the extension of the cubic-law regime in the DOS grows, and, with it, the relative density of low-frequency phonons (compared to high-frequency ones). As is well known Ziman (1960), low-frequency phonons dominate the thermal conductivity, because they have a larger mean free path and are more resilient to scattering. The fact that the overall phonon population distribution is shifted towards low-frequency phonons, thus leads to an increase of thermal conductivity with increasing confinement (decreasing thickness $L$) in this regime. We should also remark that these low-frequency phonons whose importance grows as $L$ decreases are all, practically, in-plane phonons since long wavelengths can mostly be accommodated in the $x-y$ plane of the film. Hence, this effect is going to be important for the in-plane thermal conductivity.
@@
[…]
@@
Figure 2: Comparison between the prediction of the proposed theory (solid line) and MD simulations data (circles) for the thickness-dependent thermal conductivity of ultrathin silicon films at $T=200$ K. The MD data are from Ref. Fu et al. (2020). Because the Debye-Peierls model used in the text is, of course, not fully quantitative, and the MD data from Ref. Fu et al. (2020) are also based on strong assumptions, the (thickness-independent) prefactors of $C_{v}(\omega)$ for $\omega<\omega_{\times}$ and for $\omega>\omega_{\times}$ in Eq. (13), were adjusted, respectively, to $\frac{3\hbar^{2}}{4\pi^{3}v^{4}k_{B}T^{2}}=9.5\times 10^{-67}$ and $\frac{3\hbar^{2}}{2\pi^{2}v^{3}k_{B}T^{2}}=1.45\times 10^{-63}$ (in SI units). These two values are not too far from the nominal values, at $T=200$ K, estimated to be $\frac{3\hbar^{2}}{4\pi^{3}v^{4}k_{B}T^{2}}=2.8\times 10^{-67}$ and $\frac{3\hbar^{2}}{2\pi^{2}v^{3}k_{B}T^{2}}=1.14\times 10^{-62}$, with the average speed of sound for silicon, $v\approx 7000$ m/s. All other parameters are fixed and their literature values can be found in the text.
The crucial ingredient in the above model, allowing it to reproduce the minimum in the simulated thermal conductivity of ultrathin silicon, is the cubic DOS at low frequency and its crossing over into the Debye DOS at $\omega_{\times}=2\pi v/L$. Without this form of the DOS, we have checked that the traditional FS interface scattering mechanism is unable to produce the minimum.

### v3

# Phonon-confinement theory of thermal conductivity in ultrathin silicon films

@@ I Introduction
Theoretical frameworks based on the Boltzmann equation, including scattering of phonons that bump off the boundaries of the film, as described by the Fuchs-Sondheimer (FS) theory Fuchs (1938); Sondheimer (1952), remain invaluable to explain the increase of thermal conductivity with increasing film thickness in the range of tens to hundreds of nanometers McGaughey et al. (2011). However, recent numerical investigations have revealed new unexpected behaviors upon further shrinking the thickness, that cannot be described by traditional theories. In particular, molecular simulations have highlighted the presence of a minimum in the thermal conductivity of ultrathin silicon films at about 1-2 nm thickness Fu et al. (2020). A similar behaviour, although without a clear minimum and much shallower, has been reported for argon films Fu et al. (2020).
As of today, the existing theoretical frameworks cannot explain these observation of the minimum in thermal conductivity at 1-2 nm thickness in silicon films and the flattening behaviour in argon films. Here we apply the nanoconfinement description of phonons in ultrathin few-layer films based on recent developments Yu et al. (2022); Zaccone (2024) and derive a new fundamental framework for the thermal conductivity of ultrathin films that fully takes into account the restriction in the reciprocal space occupancy of phonons due to confinement Balandin and Wang (1998). The new theory is verified in qualitative comparison with recent molecular dynamics (MD) simulations data in terms of the predicted thickness dependence.
In particular, the origin of the minimum is traced back to a crossover in the phonon density of states from a confinement-induced cubic law in frequency to the quadratic Debye law. This effect, in turn, originates from a redistribution of phonon momentum states in reciprocal space associated with a transition from the trivial Debye-sphere homotopy group to a non-trivial manifold which belongs to the homotopy class $\mathbb{Z}$.
II Model assumptions
We should emphasize the physical assumptions of the theory to be presented below. First of all, we shall always treat the ultra-thin films as 3D nanomaterials, where the atomic vibrations occur in all three spatial directions. This is true also for ultra-thin films of 1-2 nm thickness which possess a few atomic layers also along the confinement direction. Hence, we shall not consider exactly or ”ideal” 2D systems which are monolayers where the atoms vibrate only in the plane, since their thickness is much lower than the thickness considered in this paper. Indeed, for a Si monolayer, the thickness ranges from just 0.4-0.5 Å in silicene to about 3-4 Å for Si(111) monolayers Lee et al. (2013); Zhao et al. (2016); Pawlak et al. (2020); Chen et al. (2015). In what follows, instead, we will focus on films with thickness larger than $\sim 1$ nm, which thus contain at least 2-3 atomic layers. These systems are of course, fully 3D, from the point of view of phonon dynamics, as atomic vibrations take place in all three spatial directions.
Furthermore, while the theory reported here is agnostic in terms of the type of vibrational excitations that contribute to the density of states, we argue that flexural modes do not change the basic form of the DOS ($\sim\omega^{3}$) dictated by nanoconfinement, as observed experimentally and in simulations for ultra-thin films Yu et al. (2022); Fu et al. (2020); Guo et al. (2023). In these systems, the characteristic square-root of frequency trend of the low-energy phonon density of states (DOS) induced by flexural modes Kosevich (2005) is never observed.
III Derivation
@@
$\kappa=\frac{1}{3}\int_{0}^{\omega_{D}}C_{v}(\omega)v(\omega)^{2}\tau(\omega)d\omega$ (1)
where $C_{v}(\omega)$ is the frequency-dependent volumetric heat capacity, $v(\omega)$ is the phonon group velocity, and $\tau(\omega)$ is the phonon scattering time. The latter receives contribution from various processes, that include: Rayleigh-type scattering due to isolated defects, anharmonic phonon-phonon scattering, boundary scattering etc. As we are going to specialize on a simple monoatomic system such as silicon, we can use the original Debye-Peierls model without accounting for optical phonons explicitly. Then, the spectral heat capacity is given by the energy of a mode with frequency $\omega$: $U=g(\omega)\langle n(\omega)\rangle\hbar\omega$, where $g(\omega)$ is the phonon density of states (DOS), and $\langle n(\omega)\rangle=\frac{\hbar\omega}{e^{\hbar\omega/k_{B}T}-1}$ is the Bose-Einstein occupation factor. It should be noted that, using the Bose-Einstein quantum statistics, does not imply any loss of generality because at room temperature the Bose-Einstein statistics reduces to the Boltzmann statistics. Taking the derivative of the internal energy of mode $\omega$ with respect to temperature $T$, and using the Debye DOS Kittel (1996): $g(\omega)=\frac{3V}{2\pi^{2}}\frac{\omega^{2}}{v^{3}}$, the spectral heat capacity per unit volume is then:
@@
Finally, the frequency-dependent inverse of the mean free path of the phonons, which is proportional to the inverse of the average relaxation time, is given by Ward and Broido (2010); Walker and Pohl (1963); Braun et al. (2016); Callaway (1959):
$\Lambda^{-1}_{bulk}(\omega)\propto\tau^{-1}(\omega)=\frac{v}{L}+A\omega^{4}+B\omega^{2}e^{-C/T}$ (3)
where the first term on the r.h.s. is a boundary scattering term McGaughey et al. (2011); Larkin and McGaughey (2014); Braun et al. (2016) which suffices to describe size-effects in thin films with thickness $L>10$ nm. The second term on the r.h.s. represents the Rayleigh-type scattering of phonons by diluted defects and impurities (a similar quartic form is found also for more generic atomic-scale disorder Baggioli and Zaccone (2022a)). The last term on the r.h.s. is the anharmonic contribution from inelastic three-phonon (Umklapp) scattering, famously theorized by Peierls Peierls (1955) and subsequently computed by Klemens Klemens and Simon (1951).
@@
Figure 1: Left: diametral section of the corresponding geometry of $k$-space, where the Debye sphere (of radius $k_{D}$, in red) contains two symmetric ”hollow” spheres (hole pockets of radius $\pi/L$, in white) of forbidden states. These are phonon states in $k$-space that remain unoccupied due to confinement along the $z$-axis of the thin film in real space. Upon further decreasing the film thickness $L$, eventually the Debye spherical surface gets distorted, as shown in the right figure. At this point the phonon momentum states in reciprocal space are strongly redistributed, with a spreading out of states on the distorted Debye surface and many states being pushed towards the core of the red figure, resulting in a faster growth of the number of states with $k$, hence with $\omega$, explaining the higher, cubic, exponent, of the DOS derived in Yu et al. (2022). Reproduced from [R. Travaglino and A. Zaccone, Journal of Applied Physics 133, 033901 (2023)], with the permission of AIP Publishing.
However, additional effects arise when the film thickness $L$ drops below 10 nm, which are not taken into account by the above theory framework. These new effects are brought about by the spatial confinement of the quasiparticles, and can affect both thermal Zaccone (2024) and electronic transport Zaccone (2025a), as well as the superconductivity Travaglino and Zaccone (2023); Ummarino and Zaccone (2024). In a nutshell, phonons described as quantum waves have a certain wavelength. Along the confined ($z$) direction of the film, only wavelengths smaller than the thickness $L$ are allowed to exist and to populate the Debye sphere in reciprocal space. On the other hand, in the unconfined xyx-y plane, phonons with any wavelength are allowed. Along a generic direction identified by the angle $\theta$ (measured with respect to the confinement axis $z$), only phonons with wavelength:
@@
are allowed to populate the Debye sphere in $k$-space Zaccone (2025b). Phonons with larger $\lambda$ are forbidden and the corresponding points in $k$-space form two symmetric hole-pocket spheres of prohibited states within the Debye sphere, cf. Fig.1 and Figs. 1-2 in Phillips et al. (2021). As long as these two hole-pocket spheres are within the surface of the Debye sphere, the DOS is still provided by the Debye DOS quoted above. However, as the thickness $L$ is decreased, the two hole-pocket ”hollow” spheres grow up to the point that they come out of the Debye sphere (right figure in Fig. 1). At this point, the DOS is no longer given by the Debye DOS but can still be derived in simple analytical form, as follows.
@@
which shows a cubic dependence on frequency, $\sim\omega^{3}$, consistent with both experimental results (from inelastic neutron scattering) and molecular dynamics simulations reported in Ref. Yu et al. (2022). Notably, this $\omega^{3}$ behavior is observed in both crystalline and fully amorphous thin films. Furthermore, this $\omega^{3}$ law is approximately observed in MD simulations of ultrathin silicon films and ultrathin argon films Turney et al. (2010); Fu et al. (2020). This cubic DOS for ultrathin films leads to a characteristic $\sim T^{4}$ of the heat capacity with temperature at low temperatures (instead of Debye’s $\sim T^{3}$), which appears in qualitative agreement with the available data Zaccone (2024).
@@
This physical model thus explains how the confinement leads to a relative increase in the density of low-frequency vibrational modes, as indeed observed in simulations and experiments Yu et al. (2022); Turney et al. (2010); Fu et al. (2020). In particular, the excess of vibrational modes over the Debye prediction is controlled by the crossover in Eq. (10) occurring at $\omega_{\times}=\frac{2\pi}{L}v$, itself a function of the film thickness $L$. Clearly, as $L$ decreases, the extension of the cubic-law regime in the DOS grows, and, with it, the relative density of low-frequency phonons (compared to high-frequency ones). As is well known Ziman (1960), low-frequency phonons dominate the thermal conductivity, because they have a larger mean free path and are more resilient to scattering. The fact that the overall phonon population distribution is shifted towards low-frequency phonons, thus leads to an increase of thermal conductivity with increasing confinement (decreasing thickness $L$) in this regime. We should also remark that these low-frequency phonons whose importance grows as $L$ decreases are all, practically, in-plane phonons since long wavelengths can mostly be accommodated in the xyx-y plane of the film. Hence, this effect is going to be important for the in-plane thermal conductivity.
We should also emphasize that it is the relative contribution of low-frequency phonons, not their absolute count, which matters for the thermal conductivity. Ultimately, this is because thermal conductivity is a weighted integral, not a particle count. The integral sums how effectively each frequency channel transports heat, and the effectiveness is determined by the transport weight $v(\omega)^{2}\tau(\omega)$, not by the absolute number of phonons at that frequency. Hence, the relative contribution of low-frequency phonons matters because thermal conductivity is determined not by how many phonons exist, but by how strongly each frequency channel is weighted by its transport efficiency as encoded in the product $v(\omega)^{2}\tau(\omega)$ in the integrand; low-frequency acoustic modes have overwhelmingly larger efficiency, so even a small fraction of them dominates the heat flow.
@@
$+\frac{3}{2}\frac{\Lambda_{bulk}}{L}\int_{1}^{\infty}\left(\frac{1}{t^{3}}-\frac{1}{t^{5}}\right)e^{-\frac{L}{\Lambda_{bulk}}t}\,dt$ (14)
$\kappa=\frac{1}{3}\int_{0}^{\omega_{D}}C_{v}(\omega)v(\omega)S\left(\frac{\Lambda_{bulk}}{L}\right)\Lambda_{bulk}(\omega)d\omega.$ (15)
IV Physical intuition behind the transition from phonon hole-pockets to a distorted Debye surface
When an ultrathin film is confined along one spatial direction, the standing-wave condition $|k_{z}|\geq\pi/L$ suppresses all vibrational modes whose out-of-plane wavelength exceeds the film thickness $L$. In reciprocal space, this eliminates an entire set of states near $k_{z}=0$, corresponding geometrically to two spherical regions of radius $\pi/L$ that are carved out of the interior of the Debye sphere. For moderately confined films, these “hole pockets” remain entirely within the Debye sphere, and the overall structure of the allowed $k$-space volume is still close to a spherical shell with internal cavities, so that the density of states retains approximately the Debye form $g(\omega)\!\sim\!\omega^{2}$.
As the film is thinned further, the radius of the forbidden pockets grows until these cavities pierce the boundary of the Debye sphere. Beyond this point, the geometry of the allowed $k$-space region is no longer a sphere with removed interior volume: the Debye surface itself is reshaped, becoming strongly compressed along the confinement direction and expanded in the transverse directions. Phonon states that would normally populate a spherical shell are forced onto this distorted manifold, leading to a faster growth of the number of states with increasing $k$ and thereby producing a low-frequency density of states that scales as $g(\omega)\!\sim\!\omega^{3}$ instead of the Debye $\omega^{2}$ law.
This geometric transition in reciprocal space has direct consequences for heat transport. The redistribution of vibrational states toward low frequencies enhances the relative population of long-wavelength acoustic modes, which possess higher group velocities and longer mean free paths. Thus, the distortion of the Debye surface under strong confinement naturally leads to an increase in the contribution of low-frequency phonons to the thermal conductivity, providing the physical basis for the non-monotonic thickness dependence observed in ultrathin silicon films.
V Negligible contribution of optical phonons to heat transport
In crystalline semiconductors such as silicon, the lattice thermal conductivity is overwhelmingly dominated by acoustic phonons. This can be understood by considering several factors.
@@ V Negligible contribution of optical phonons to heat transport
Second, while optical phonons lie at higher energies (e.g. $\sim 60\penalty 10000\ \text{meV}$ in Si) and are only partially populated at room temperature, acoustic phonons are numerous and thermally excited across the full spectrum.
Third, optical phonons experience short lifetimes due to strong anharmonic decay, whereas acoustic phonons—particularly the long-wavelength modes—exhibit longer mean free paths and hence contribute far more effectively to heat transport. Quantitative branch-resolved Boltzmann transport calculations confirm this picture: at $300\penalty 10000\ \text{K}$, more than $95\%$ of the thermal conductivity of silicon arises from the three acoustic branches (LA $\sim 35\%$, $_{1}\sim 28\%, TA$$_{2}\sim 32\%$), while all optical branches together account for only about $5\%$ of the total Wang and Huang (2014).
For all these reasons, it is appropriate to focus the theoretical description on the role of acoustic phonons, since that is also the range where confinement effects are more important Fu et al. (2020), and use an averaged Debye-type schematic description for the optical sector, as done above.
VI Additional phonon scattering mechanisms in ultrathin films
Beyond the intrinsic phonon–phonon and diffuse boundary scattering processes discussed above, a comprehensive description of thermal transport in experimentally realized ultrathin films requires incorporating several additional scattering channels that become increasingly important at the nanoscale. First, interface roughness scattering represents a major source of momentum relaxation in thin films grown on substrates or fabricated via vapor-phase deposition. Atomic-scale height fluctuations at free or buried interfaces lead to wavelength-dependent partial specularity, with the specularity parameter often modeled as $p(\lambda)\simeq\exp[-16\pi^{3}\eta^{2}/\lambda^{2}]$, where $\eta$ denotes the root-mean-square roughness Ziman (1960); Soffer (1967). Such roughness strongly suppresses cross-plane heat flow, as short-wavelength phonons experience nearly fully diffuse scattering, while only long-wavelength acoustic modes retain partial specularity.
Second, material defects such as substitutional impurities, dopants, vacancies, mass-disorder inhomogeneities, as well as structural disorder, introduce Rayleigh-type elastic scattering, which produces a characteristic $\omega^{4}$ dependence of the inverse relaxation time Klemens (1955a); Carruthers (1961); Baggioli and Zaccone (2022b). In doped silicon, impurity scattering can dominate over anharmonic Umklapp processes across a wide frequency range, substantially reducing the population of mid- and high-frequency phonons. Line defects such as dislocations also contribute significantly: the long-range strain fields surrounding dislocations scatter phonons via a relaxation time that depends sensitively on branch and frequency Klemens (1955b); Granato and Lücke (1966). Grain boundaries, common in polycrystalline or nanocrystalline films, act as partially transmitting interfaces whose reflection probability depends on phonon wavelength, incident angle, and polarization; they therefore constitute an additional source of diffuse scattering and mode conversion Mayadas and Shatzkes (1969).
Finally, in supported films or multilayers, substrate and interfacial coupling provide further channels for phonon momentum relaxation. Acoustic mismatch, partial phonon transmission into the substrate, and the presence of interfacial layers (e.g. native oxides) reduce the effective mean free path of phonons impinging on the interface Swartz and Pohl (1989). These mechanisms alter the boundary conditions experienced by the confined modes and modify the balance between in-plane and cross-plane heat transport. Taken together, these extrinsic scattering mechanisms—interface roughness, mass disorder, point and line defects, grain boundaries, and coupling to the substrate—play an essential role in determining the thermal conductivity of ultrathin films in realistic experimental settings and must therefore be considered alongside the intrinsic phonon-confinement effects described in this work.
In the model presented here, effects due to disorder mentioned above are effectively taken into account by the Rayleigh-type $\omega^{4}$ contribution to the phonon scattering. The other effects discussed above can be easily integrated within the proposed model for future applications thereof to experimental data.
VII Comparison with simulations data and discussion
@@
Figure 2: Comparison between the prediction of the proposed theory (solid line) and MD simulations data (circles) for the thickness-dependent thermal conductivity of ultrathin silicon films at $T=200$ K. The MD data are from Ref. Fu et al. (2020). Because the Debye-Peierls model used in the text is, of course, not fully quantitative, and the MD data from Ref. Fu et al. (2020) are also based on strong assumptions, the (thickness-independent) prefactors of $C_{v}(\omega)$ for $\omega<\omega_{\times}$ and for $\omega>\omega_{\times}$ in Eq. (13), were adjusted, respectively, to $\frac{3\hbar^{2}}{4\pi^{3}v^{4}k_{B}T^{2}}=9.5\times 10^{-67}$ and $\frac{3\hbar^{2}}{2\pi^{2}v^{3}k_{B}T^{2}}=1.45\times 10^{-63}$ (in SI units). These two values are not too far from the nominal values, at $T=200$ K, estimated to be $\frac{3\hbar^{2}}{4\pi^{3}v^{4}k_{B}T^{2}}=2.8\times 10^{-67}$ and $\frac{3\hbar^{2}}{2\pi^{2}v^{3}k_{B}T^{2}}=1.14\times 10^{-62}$, with the average speed of sound for silicon, $v\approx 7000$ m/s. All other parameters are fixed and their literature values can be found in the text. The MD simulations data are normalized by the bulk value of the MD simulations, 414 W/m-K, while the theory line is normalized by an overall factor 9.2.
The crucial ingredient in the above model, allowing it to reproduce the minimum also observed in the simulated thermal conductivity of ultrathin silicon, is the cubic DOS at low frequency and its crossing over into the Debye DOS at $\omega_{\times}=2\pi v/L$. Without this form of the DOS, we have checked that the traditional FS interface scattering mechanism is unable to produce the minimum.
@@
While different assumptions are used in the simulations of Ref. Fu et al. (2020) and in the present analytical theory, the presence of the minimum appears to be a robust feature, as well as the theoretical prediction of thermal conductivity increasing with film thickness for films thicker than 2 nm, as observed in many different materials Park et al. (2025).
VIII Conclusions